OpenStudy (anonymous):

Find the value of k so that the line passing through the given points has slope m. Write an equation of the line in point-slope form. (k,4k), (k+2,3k), m=-1

OpenStudy (aum):

\[ \text{Slope = }\frac{y_2-y_1}{x_2-x_1} = \frac{3k-4k}{(k+2)-k} = \frac{-k}{2} = -1\\ -k = -2 \\ k = 2 \\ \text{Points are: }(2,8); ~~(4,6) \\ \text{Equation in point-slope form is:} \\ y - y_1 = m(x - x_1) \\ y - 8 = -1(x - 2) \]